# Quasinormal modes in semi-open systems

### v2

@@ Abstract
Astrophysical black holes are open systems which, when perturbed, radiate quasi-normal modes (QNMs) to infinity. By contrast, laboratory analogues are necessarily finite-sized, presenting a potential obstacle to exciting QNMs in the lab. In this study, we investigate how the QNM spectrum of a toy-model black hole is modified when the system is enclosed by a partially reflecting wall. Counter to expectation, we demonstrate that QNMs not only persist in finite-sized systems, but the number of detectable modes increases. Furthermore, we show that QNMs in this set-up can be easily excited by incoherent background noise. Our findings align with studies exploring the spectral stability of black holes, such as those examining small modifications of the surrounding gravitational field. Importantly, our work paves the way for exploring spectral stability in laboratory systems, enabling experimental investigation of black hole features in closed systems.
@@
Introduction.—The success of the black hole spectroscopy program is founded on the use of gravitational waves produced by merging binaries as a probe of the system’s properties [1]. At late times during the so-called ringdown phase, the remnant emits a characteristic signal as it sheds energy and relaxes into its final state. This signal consists of a discrete set of damped harmonic oscillators referred to as quasinormal modes (QNMs) [2, 3, 4, 5, 6], which are natural resonances of linear open systems. Next generation gravitational wave detectors such as LISA and Einstein telescope will profit from enhanced sensitivity [7, 8], making it paramount to understand how the ringdown is influenced by various high-order effects, e.g. environmental factors [9] and nonlinearities [10].
@@
Recent investigation into the spectral stability of black holes has demonstrated the susceptibility of the QNM spectrum to small, local perturbations to the scattering potential experienced by gravitational waves around a black hole [11, 12, 13, 14]. In some cases, the spectrum can even admit additional long-lived modes called quasi-bound states, potentially leading to higher-order effects as a result of their long lifetimes [15, 16]. Understanding such modifications of the QNM spectrum may have important consequences for the black hole spectroscopy programme [17, 18]. In this regard, it is natural to consider non-perturbative scenarios where energy is retained by the system due to an effective reflection mechanism acting on escaping radiation. Such situations arise naturally in gravity in the context of neutron stars [4], exotic compact objects [19, 20, 21, 22] and asymptotically anti-de Sitter spacetimes [6].
@@
Gravity simulators present an ideal platform where such an enquiry can be pursued. These are physical systems in which excitations behave as though they propagate on a curved spacetime, permitting the experimental investigation of elusive phenomena like Hawking radiation [23, 24, 25, 26] and rotational superradiance [27, 28, 29, 30, 31]. Notably, ringdown signals have been studied in polariton superfluids [32], optical solitons [33] and in hydrodynamic systems [34, 35]. Although treating these platforms as effectively open systems is usually justified by a combination of energy dissipation, limited temporal evolution and engineering of absorptive boundaries, a complete analysis should not neglect finite-size effects resulting from confinement mechanisms. This is particularly true of our recent superfluid helium gravity simulator [36], where dissipation is strongly reduced and many reflections from the container wall can occur during the experimental timeframe.
@@
All of this serves to emphasise the importance of understanding how finite-size effects modify the QNM spectrum in the context of both black hole physics and gravity simulators. In this Letter, we study the effect of a partially reflective boundary on the QNM spectrum of the Pöschl-Teller (PT) potential, a model which appears in both black hole physics [6, 2] and analogue systems [37, 33], demonstrating how finite-size effects endow the system with an additional control parameter that can be exploited in the experimental study of QNMs.
@@
General set-up.—A simple system which captures the essential physics of QNMs is the one-dimensional wave equation with a repulsive potential $V(x)$ [4, 6],
@@
In this case, (1) has an analytical solutions in the form $e^{-i\omega t}\psi(x)$. Defining a new variable $y=1/(1+e^{2\alpha x})$, the spatial part can be written as,
@@
with the $+$ (--) sign for $a$ ($b$). Without loss of generality, from now on we take $V_{0}=1$. To impose the resonance condition (5), we need to define the reflection coefficient. Assuming that $V(x_{b})$ is close to vanishing, (Quasinormal modes in semi-open systems) is well-approximated by its asymptotic expansion and the in- and out-going wave contributions can be identified (see Appendix B). Their ratio gives the reflection coefficient,
@@
Figure 1: a. Trajectories in the complex plane of the first 11 resonances of the PT potential with a partially-reflective barrier relative to the fundamental QNM $\omega_{n=0}^{\varepsilon=0}\equiv\tilde{\omega}_{r}+i\tilde{\omega}_{i}$. Here we take $\alpha=0.4$ and the position of the wall is fixed at $x_{b}=10$. The large yellow dots represent QNMs of the open system, whereas the blue lines are obtained varying $\varepsilon$. Coloured markers label the resonances at particular values of reflectivity. b. Trajectories of the resonances as reflectivity increases, for $\alpha=0.5$ (Green, dashed), $=0.4$ (Blue, solid), $=0.3$ (Orange, dotted) and $x_{b}=10$. The yellow dots are $\omega_{n=0,1}^{\varepsilon=0}$ and the lines end at $\varepsilon=1$. c. The migration of the resonance is compared against the vector field $\mathrm{H}(\omega)$ for $\alpha=0.4$. The red cross is placed on the repelling point, i.e. the pole of $\mathrm{H}$.
Resonances of the semi-open system are obtained by numerically solving (5). In Fig. 1a, we display their continuous migration in the complex plane as $\varepsilon$ is varied. Large yellow dots represent QNMs of the open system for $\varepsilon=0$ and, as $\varepsilon$ is increased, the resonances migrate down toward the real line. We can identify three qualitatively different types of QNM when $\varepsilon=1$. Modes with $\mathrm{Re}[\omega]=0$ correspond to purely damped solutions (discussed in [40]). Modes with $0<\mathrm{Re}[\omega]<\tilde{\omega}_{r}$ are damped oscillatory modes which are localised between the hard wall and the barrier $V$. These are quasibound states (QBS) which need to tunnel through $V$ to escape to $x\to-\infty$, leading to long lifetimes. Finally, modes with $\mathrm{Re}[\omega]>\tilde{\omega}_{r}$ pass over the barrier and experience only a small amount of reflection due to the inhomogeneous nature of $V$. Consequently, they are more strongly damped and can be identified with the new overtones of the system. These overtones possess distinctly different real frequencies and an imaginary part which is reduced significantly relative to the open system.
@@
Numerical simulations.—A consequence of the reflective boundary is that the new overtones have distinct real frequencies and longer lifetimes, facilitating their detection. We demonstrate this via a numerical simulation. The wave equation (1) is simulated using a method of lines algorithm (details in Appendix A). In Fig. 2, a Gaussian pulse is evolved toward the barrier from the side of the open boundary. We illustrate the response signal, in both the time and frequency domains (left and right panels respectively) after the initial pulse has passed for various $\varepsilon$, i.e. as the boundary is made more reflective, focusing on values which make the decay times of $n=0,1$ comparable (see Fig. 1). For the open system, we observe the characteristic QNM waveform, corresponding to a singly peaked frequency spectrum. However, for semi-open systems, the signal becomes more complicated, characterised by interference between multiple frequencies in the time domain which correspond to more than one peak in Fourier space. Notably, the peak corresponding to $\omega^{\varepsilon=0}_{n=0}$ splits in two, the left peak corresponding to the highest frequency QBS and the right peak to the first of the new overtones. For a partially absorbing wall, we observe that this overtone is actually the dominant signal when we measure near the open boundary. When the wall is fully reflective, the dominant signal comes from the QBS due to their long lifetimes, although our simulations suggest that the first overtone is still detectable. In Fig. 2c, we illustrate the full spatial waveform as a function of frequency, illustrating how the different resonant modes (indicated by bright horizontal lines) are bound by the potential barrier. As expected, modes scattering with the wider part of the barrier experience less leakage. Note that although format in Fig. 2a will be familiar to gravitational community, the equivalent representations in panels b and c emerge naturally in the data analysis of analogue experiments [35, 42] and carry more information about the spatial profile of the modes.
@@
Although Gaussian pulses are frequently used in numerical studies to stimulate the system [6], a more natural perturbing mechanism from the experimental perspective is broadband mechanical noise [35, 36]. In Fig. 3, we take $\Psi(x,t=0)=0$ as initial condition but generate random noise on the side of the open boundary at each time step for the duration of the simulation. We then Fourier transform the signal after the noise has propagated across the system and take the average on the side of the barrier near the wall. We collect statistics over 500 iterations and display the mean and the variance. In the open system, there is no distinct signal; at high frequency we observe the noise level and at lower frequencies a suppression due to the presence of the barrier. For partial absorption, we observe a series of peaks emerging above the noise level, which become very pronounced for the fully reflective wall. Notably, the first overtone appears as a distinct signal in the data, suggesting that the combination of a finite size system with mechanical noise may be sufficient ingredients to detect QNM overtones in analogue experiments. This method of resonant stimulation is akin to that of the driven, damped harmonic oscillator, where the energy input from the driving force is balanced by the energy loss due to dissipation (in our case resulting from the open boundary).
@@
Discussion.—Finite-size constraints of analogue experiments might initially appear as an obstacle to simulating black hole ringdown - a phenomenon typically associated with open (i.e. spatially unbounded) systems. Counter to this expectation, our results demonstrate that the introduction of such confining mechanisms, in fact, becomes a valuable tool for studying the quasinormal modes (QNMs) and the role of the overtones in shaping the ringdown signal. The set-up we have presented (based on a Poschl-Teller potential) captures the essential physics of experimentally realisable black hole analogues, where the partially reflective boundary models the experimental confinement and the open boundary represents the analogue black hole horizon. As the reflectivity of the closed boundary is increased from zero, the QNMs of the open system migrate continuously along trajectories in the complex frequency plane toward the real axis. In addition to the familiar quasibound states of semi-open systems, the modified spectrum contains a hierarchy of overtones with distinct oscillation frequencies and substantially longer lifetimes compared to the open system.
@@
These two features significantly increase the prospect of a detecting QNM overtones in an analogue experiment, as exemplified by our numerical simulations. An added consequence is that the frequency response of the semi-open system to a random driving signal is sharply peaked around the QNMs, a feature which is absent in the open system as it relies on constructive interference with reflected waves. A measurement of such a signal could prove particularly natural for experiments where noise (e.g. thermal or mechanical) is always present. Indeed, this mechanism may be responsible for the QNM oscillations measured in [37].
@@
From a gravity perspective, the semi-open problem has been addressed in the context of neutron stars/compact objects and anti-de Sitter spacetimes. The focus (in the case of the former) is usually on the new bound states that arise in the spectrum, and (in the latter) on echoes of the usual QNM response. In our case, where the reflecting boundary is close to the peak of the scattering potential, we have shown that the new overtones are a prominent feature of linear response, and that random noise is a viable probe of the spectrum. In addition to the aforementioned astrophysical settings, stimulating QNMs in this fashion may also be relevant for a black hole surrounded by matter [43, 44, 9] or hypothetical ultra-light fields [45, 46].
Furthermore, discerning the roles of nonlinearities and overtones in shaping the ringdown waveform is a central topic of research in the gravitational wave community [47, 48]. Understanding the interplay of these two competing effects is therefore desirable. Finite size black hole simulators may be useful in this regard, since both the nonlinearities and overtone lifetimes can be experimentally controlled (the former through signal amplitudes and the latter using the boundary reflectivity). This feature stands out as an enhanced capability of semi-open systems compared with gravitational black holes which radiate to infinity, where a consequence of the no-hair theorem is that the overtone lifetime is not an adjustable parameter [6].
@@
Acknowledgements.—L.S. and S.P. would like to thank Théo Torres for providing valuable insight on spectral stability based on unpublished work with Sam Dolan. We are grateful to Patrik Švančara, Pietro Smaniotto and Thomas Sotiriou for helpful discussions and comments on the manuscript. We also thank Paolo Pani for feedback. L.S. and S.W. gratefully acknowledge the support of the Leverhulme Research Leadership Award (RL-2019-020). S.P., R.G. and S.W. extend their appreciation to the Science and Technology Facilities Council for their generous support in Quantum Simulators for Fundamental Physics (QSimFP), (ST/T005858/1 and ST/T006900/1), as part of the UKRI Quantum Technologies for Fundamental Physics program. S.W. also acknowledges the Royal Society University Research Fellowship (UF120112). R.G. acknowledges support from the Perimeter Institute. Research at Perimeter Institute is supported by the Government of Canada through the Department of Innovation, Science and Economic Development Canada and by the Province of Ontario through the Ministry of Research, Innovation and Science.
Appendix A Numerical simulations
$\partial_{x}^{2}\Psi_{k}=\frac{\Psi_{k+1}-2\Psi_{k}+\Psi_{k-1}}{h^{2}}.$ (12)
We then define $\Pi=\partial_{t}\Psi$ and cast (1) in the form,
$\partial_{t}\begin{pmatrix}\Psi\\ \Pi\end{pmatrix}=\begin{pmatrix}0&1\\ \partial_{x}^{2}-V&0\end{pmatrix}\begin{pmatrix}\Psi\\ \Pi\end{pmatrix},$ (13)
$\partial_{x}\Psi_{k}=\frac{\Psi_{k+1}-\Psi_{k-1}}{2h},$ (14)
$\partial_{x}^{2}\Psi_{1}=\frac{2\Psi_{2}-2\Psi_{k-1}}{h^{2}}-\frac{2\Pi_{1}}{h}.$ (15)
$\partial_{t}\begin{pmatrix}\Psi\\ \Pi\end{pmatrix}=\begin{pmatrix}\mathbf{0}&\mathbf{I}\\ {}^{\mathrm{N}}\partial_{x}^{2}-V&B\end{pmatrix}\begin{pmatrix}\Psi\\ \Pi\end{pmatrix},$ (16)
where $B=-(2/h)\mathrm{diag}(\mu,0,...0,1)$ is a diagonal matrix with non-zero entries in the top-left and bottom-right corner. The parameter $\mu$ allows us to tune between a fully reflective Neumann boundary condition on one side for $\mu=0$ and a completely open boundary for $\mu=1$. The relation between $\mu$ and the reflectivity $\varepsilon$ used in the main text is,
$\mu=\frac{1-\varepsilon}{1+\varepsilon}.$ (17)
Appendix B Pöschl-Teller resonance condition

### v3

# On the origin of quasinormal modes in semi-open systems

@@ Abstract
Astrophysical black holes are open systems which, when perturbed, radiate quasi-normal modes (QNMs) to infinity. By contrast, laboratory analogues are necessarily finite-sized, presenting a potential obstacle to exciting QNMs in experiments. We explore how the QNM spectrum of a toy-model black hole changes when enclosed by a partially reflecting wall with adjustable reflectivity. Our results reveal a continuous connection between the QNM spectra of open and finite-sized systems. Additionally, we demonstrate that QNMs in this setup are easily excited by incoherent background noise. This work opens new avenues for studying QNMs of black holes and compact objects in laboratory settings, where finite-size effects and noise are unavoidable.
@@
Introduction.—The success of the black hole spectroscopy program is founded on the use of gravitational waves produced by merging binaries as a probe of the system’s properties [1]. At late times during the so-called ringdown phase, the remnant emits a characteristic signal as it sheds energy and relaxes into its final state. This signal consists of a discrete set of damped harmonic oscillators referred to as quasinormal modes (QNMs) [2, 3, 4, 5, 6, 7], which are natural resonances of linear open systems. Next generation gravitational wave detectors such as LISA and Einstein telescope will profit from enhanced sensitivity [8, 9], making it paramount to understand how the ringdown is influenced by various high-order effects, e.g. environmental factors [10] and nonlinearities [11].
@@
Recent investigation into the spectral stability of black holes has demonstrated the susceptibility of the QNM spectrum to small, local perturbations to the scattering potential experienced by gravitational waves around a black hole [12, 13, 14, 15]. In some cases, the spectrum can even admit additional long-lived modes called quasi-bound states, potentially leading to higher-order effects as a result of their long lifetimes [16, 17]. Understanding such modifications of the QNM spectrum may have important consequences for the black hole spectroscopy programme [18, 19]. In this regard, it is natural to consider non-perturbative scenarios where energy is retained by the system due to an effective reflection mechanism acting on escaping radiation. Such situations arise naturally in gravity in the context of neutron stars [4], exotic compact objects [20, 21, 22, 23] and asymptotically anti-de Sitter spacetimes [6, 7].
@@
Gravity simulators present an ideal platform where such an enquiry can be pursued. These are physical systems in which excitations behave as though they propagate on a curved spacetime, permitting the experimental investigation of elusive phenomena like Hawking radiation [24, 25, 26] and rotational superradiance [27, 28, 29, 30, 31]. Notably, ringdown signals have been studied in polariton superfluids [32], optical solitons [33] and in hydrodynamic systems [34, 35]. Although treating these platforms as effectively open systems is usually justified by a combination of energy dissipation, limited temporal evolution and engineering of absorptive boundaries, a complete analysis should not neglect finite-size effects resulting from experimental confinement mechanisms, e.g. in classical fluids [35], superfluids [36], cold-atomic-gases [37], polariton fluids [38] etc.
@@
All of this serves to emphasise the importance of understanding how finite-size effects modify the QNM spectrum. In this Letter, we study the effect of a partially reflective boundary on the QNM spectrum of the Pöschl-Teller (PT) potential. This system is sufficiently versatile to capture qualitative features of potentials around astrophysical black holes and their analogues [6, 7, 2, 39, 33] whilst remaining analytically solvable. We demonstrate that finite-size effects endow the system with an additional control parameter that can facilitate the experimental detection of QNMs, enabling control of their lifetimes and their excitation by experimental noise.
@@
General set-up.—A simple system which captures the essential physics of QNMs is the one-dimensional wave equation with a repulsive potential $V(x)$ [4, 6, 7],
@@
In this case, (1) has an analytical solutions in the form $e^{-i\omega t}\psi(x)$. Defining a new variable $y=1/(1+e^{-2\alpha x})$, the spatial part can be written as,
@@
with the $+$ (--) sign for $a$ ($b$). Without loss of generality, from now on we take $V_{0}=1$, so that the problem is characterised by the two dimensionless scales $\alpha$ and $x_{b}$. To impose the resonance condition (5), we need to define the reflection coefficient. Assuming that $V(x_{b})$ is close to vanishing, (On the origin of quasinormal modes in semi-open systems) is well-approximated by its asymptotic expansion and the in- and out-going wave contributions can be identified (see Appendix A). Their ratio gives the reflection coefficient,
@@
Figure 1: a. Trajectories in the complex plane of the first 11 resonances of the PT potential with a partially-reflective barrier relative to the fundamental QNM $\omega_{n=0}^{\varepsilon=0}\equiv\tilde{\omega}_{r}+i\tilde{\omega}_{i}$. Here we take $\alpha=0.4$ and the position of the wall is fixed at $x_{b}=10$. The large yellow dots represent QNMs of the open system, whereas the blue lines are obtained varying $\varepsilon$. Coloured markers label the resonances at particular values of reflectivity. b. Trajectories of the resonances as reflectivity increases, for $\alpha=0.3$ (Green, dashed), $=0.4$ (Blue, solid), $=0.5$ (Orange, dotted) and $x_{b}=10$. The yellow dots are $\omega_{n=0,1}^{\varepsilon=0}$ and the lines end at $\varepsilon=1$. c. The migration of the resonance is compared against the vector field $\mathrm{H}(\omega)$ for $\alpha=0.4$. The red cross is placed on the repelling point, i.e. the pole of $\mathrm{H}$.
Resonances of the semi-open system are obtained by numerically solving (5). In Fig. 1a, we display their continuous migration in the complex plane as $\varepsilon$ is varied, fixing $x_{b}=10$ and $\alpha=0.4$. The trajectories are presented relative to the fundamental QNM $\omega_{n=0}^{\varepsilon=0}\equiv\tilde{\omega}_{r}+i\tilde{\omega}_{i}$. Large yellow dots represent QNMs of the open system for $\varepsilon=0$ and, as $\varepsilon$ is increased, the resonances migrate down toward the real line. We can identify three qualitatively different types of QNM when $\varepsilon=1$. Modes with $\mathrm{Re}[\omega]=0$ correspond to purely damped solutions (discussed in [42]). Modes with $0<\mathrm{Re}[\omega]<\tilde{\omega}_{r}$ are damped oscillatory modes which are localised between the hard wall and the barrier $V$. These are quasibound states (QBS) which need to tunnel through $V$ to escape to $x\to-\infty$, leading to long lifetimes. Finally, modes with $\mathrm{Re}[\omega]>\tilde{\omega}_{r}$ pass over the barrier and experience only a small amount of reflection due to the inhomogeneous nature of $V$. Consequently, they are more strongly damped and can be identified with the new overtones of the system. These overtones possess distinctly different real frequencies and an imaginary part which is reduced significantly relative to the open system. Note, we are mainly interested in the regime of small $\alpha$ (wide potentials) where the decay rate of $\omega^{\varepsilon=0}_{n=0}$ is an order of magnitude smaller than its oscillation frequency, which is the astrophysically relevant case [44]. The scale $x_{b}$ then determines the number of QBS in semi-open system.
@@
Numerical simulations.—A consequence of the reflective boundary is that the new overtones have distinct real frequencies and longer lifetimes, facilitating their detection. We demonstrate this via a numerical simulation. The wave equation (1) is simulated using a method of lines algorithm. Spatial derivatives are implemented using 3-point finite-difference stencils and the evolution is performed using a fourth-order Runge-Kutta algorithm, whilst the permeable wall is implemented by specifying a relation between the $x$ and $t$ derivatives at one of the boundaries (see Appendix B for details). In Fig. 2, a Gaussian pulse $\exp(ik_{0}x-(x-x_{0})^{2}/2\sigma^{2})$, with $k_{0}=2$, centred at $x_{0}=-7.5$ (near the open boundary), and $\sigma^{2}=5$ is evolved toward the barrier. We illustrate the response signal, in both the time and frequency domains (left and right panels respectively) after the initial pulse has passed for various $\varepsilon$, i.e. as the boundary is made more reflective, focusing on values which make the decay times of $n=0,1$ comparable (see Fig. 1). For the open system, we observe the characteristic QNM waveform, corresponding to a singly peaked frequency spectrum. However, for semi-open systems, the signal becomes more complicated, characterised by interference between multiple frequencies in the time domain which correspond to more than one peak in Fourier space. Notably, the peak corresponding to $\omega^{\varepsilon=0}_{n=0}$ splits in two, the left peak corresponding to the highest frequency QBS and the right peak to the first of the new overtones. For a partially absorbing wall, we observe that this overtone is actually the dominant signal when we measure near the open boundary. When the wall is fully reflective, the dominant signal comes from the QBS due to their long lifetimes, although our simulations suggest that the first overtone is still detectable. In Fig. 2c, we illustrate the full spatial waveform as a function of frequency, illustrating how the different resonant modes (indicated by bright horizontal lines) are bound by the potential barrier. As expected, modes scattering with the wider part of the barrier experience less leakage. Note that although format in Fig. 2a will be familiar to gravitational community, the equivalent representations in panels b and c emerge naturally in the data analysis of analogue experiments [35, 45] and carry more information about the spatial profile of the modes.
@@
Although Gaussian pulses are frequently used in numerical studies to stimulate the system [6], a more natural perturbing mechanism from the experimental perspective is broadband mechanical noise [35, 36]. In Fig. 3, we take $\Psi(x,t=0)=0$ as the initial condition but generate random noise on the side of the open boundary at each time step for the duration of the simulation. We then Fourier transform the signal after the noise has propagated across the system and take the average on the side of the barrier near the wall. We collect statistics over 500 iterations and display the mean and the variance. In the open system, there is no distinct signal; at high frequency we observe the noise level and at lower frequencies a suppression due to the presence of the barrier. For partial absorption, we observe a series of peaks emerging above the noise level, which become very pronounced for the fully reflective wall. Notably, the first overtone appears as a distinct signal in the data, suggesting that the combination of a finite size system with mechanical noise may be sufficient ingredients to detect QNM overtones in analogue experiments. This method of resonant stimulation is akin to that of the driven, damped harmonic oscillator, where the energy input from the driving force is balanced by the energy loss due to dissipation (in our case resulting from the open boundary).
@@
Discussion.—Finite-size constraints of analogue experiments might initially appear as an obstacle to simulating black hole ringdown – a phenomenon typically associated with open (i.e. spatially unbounded) systems. Counter to this expectation, our results demonstrate that the introduction of such confining mechanisms, in fact, becomes a valuable tool for studying the quasinormal modes (QNMs) and the role of the overtones in shaping the ringdown signal. The set-up we have presented (based on a Pöschl-Teller potential) captures the essential physics of experimentally realisable black hole analogues, where the partially reflective boundary models the experimental confinement and the open boundary represents the analogue black hole horizon. As the reflectivity of the closed boundary is increased from zero, the QNMs of the open system migrate continuously along trajectories in the complex frequency plane toward the real axis. In addition to the familiar quasibound states of semi-open systems, the modified spectrum contains a hierarchy of overtones with distinct oscillation frequencies and substantially longer lifetimes compared to the open system. Qualitatively similar results are expected for any asymptotically flat scattering potential with a single peak, e.g. as encountered around Schwarzschild black holes [46].
@@
These two features significantly increase the prospect of detecting QNM overtones in an analogue experiment, as exemplified by our numerical simulations. An added consequence is that the frequency response of the semi-open system to a random driving signal is sharply peaked around the QNMs, a feature which is absent in the open system as it relies on constructive interference with reflected waves. A measurement of such a signal would be particularly natural for experiments where noise (e.g. thermal or mechanical) is always present. Indeed, we recently performed experiments on a confined draining vortex in superfluid 4He [47], using the experimental noise to stimulate resonant spectral peaks. This mechanism may also be responsible for the QNM oscillations measured in classical fluid experiments [39].
@@
Besides their astrophysical relevance, QNMs are important in a wide variety of subfields, including non-Hermitian systems [48], strongly coupled field theories [49, 50], optical cavities [51, 52], semiclassical gravity [53, 54] and quantum vortex dynamics [55, 56]. By connecting QNMs of the fully open system with those of the semi-open one, we anticipate our results will inspire new crossovers between black hole perturbation theory and the areas mentioned above.
From a gravity perspective, the semi-open problem has been addressed in the context of anti-de Sitter spacetimes, compact objects and neutron stars. Long-lived modes are known to develop in all these cases [6, 23, 57], with the QBS and new overtones in our model mirroring w-modes of neutron stars called trapped and curvature modes. A finite-sized analogue system with variable reflectivity could therefore be used to experimentally investigate the transition between the neutron star regime (full reflection) and the compact object regime (partial reflection), using mechanical noise in the set-up as a probe of the spectrum. Stimulating QNMs with noise may also be relevant in purely astrophysical settings, such as black holes surrounded by matter [58, 59, 10] or hypothetical ultra-light fields [60, 61].
Discerning the roles of nonlinearities and overtones in shaping the ringdown waveform is a key topic in gravitational wave research [62, 63], and understanding their interplay is highly desirable. Finite-sized black hole simulators provide a valuable platform for this, allowing experimental control over both nonlinearities (via signal amplitudes) and overtone lifetimes (through boundary reflectivity)–a capability absent in gravitational black holes, where the no-hair theorem fixes overtone lifetimes [6]. Moreover, such analogue systems enable the study of complex nonlinear interactions relevant to black hole physics, including QCD-axions [64] and photon-plasma effects [65]. While previous studies have explored basic nonlinear terms, e.g., quartic self-couplings [66], gravity simulators offer the potential to probe higher-order interactions experimentally, a direction we will pursue in future work.
We conclude by emphasising that typical methods of stimulating QNMs with wave packets fail in confined systems, since reflections from the boundary obscure the signal, as revealed by earlier experiments [35]. Furthermore, since real experiments are intrinsically messy and overtones are short-lived, resolving them above the noise can prove difficult. Whilst this would suggest that finite-size effects and experimental noise are obstacles to the experimental study of QNMs, our results demonstrate that these features can actually be used to the experimentalist’s advantage.
@@
Acknowledgements.—L.S. and S.P. would like to thank Théo Torres for providing valuable insight on spectral stability based on unpublished work with Sam Dolan. We are grateful to Patrik Švančara, Pietro Smaniotto and Thomas Sotiriou for helpful discussions and comments on the manuscript. We also thank Paolo Pani, Kostas Kokkotas and Sebastian Völkel for feedback. L.S. and S.W. gratefully acknowledge the support of the Leverhulme Research Leadership Award (RL-2019-020). S.P., R.G. and S.W. extend their appreciation to the Science and Technology Facilities Council for their generous support in Quantum Simulators for Fundamental Physics (QSimFP), (ST/T005858/1, ST/Y004450/1, and ST/T006900/1), as part of the UKRI Quantum Technologies for Fundamental Physics program. S.W. also acknowledges the Royal Society University Research Fellowship (UF120112). R.G. acknowledges support from the Perimeter Institute. Research at Perimeter Institute is supported by the Government of Canada through the Department of Innovation, Science and Economic Development Canada and by the Province of Ontario through the Ministry of Research, Innovation and Science.
Supplemental Material
Appendix A Pöschl-Teller resonance condition